# A Survey of Secure Semantic Communications

### v1

@@ 4.3 Defensive Methods of Backdoor Attacks
Figure 9: Classification of neural network backdoor defense [131].
@@ 4.3.1 Data Processing
    Clustering based Defense. For contaminated datasets, toxic data can be separated from normal data according to cluster analysis. Activation Clustering is a defense method proposed by Chen [134] et al. The key idea is to analyze the activation mode of network for the training data, to determine whether the data is poisoned and which data points are toxic. For each sample in the training set, AC records the activation output of the hidden layer. These activation outputs are flattened into one-dimensional vectors and classified based on their labels. After dimensionality reduction, these vectors are clustered and analyzed into toxic and normal patterns. The data is analyzed by exclusionary reclassification, silhouette score, size comparison and many other ways to filter out the toxic data. Based on the labeled toxic data, the system can further train the model until convergence to achieve the purpose of backdoor repairing. One more clustering approach is Heatmap Clustering proposed by Schulth et al. [135] This method utilizes Explainable Artificial Intelligence techniques, specifically Layer-wise Relevance Propagation, to generate heatmaps. The heatmap shows the correlation between each pixel point of the input sample and the network predictions, which can be visualized as a two-dimensional array where the color shades indicate the strength of the relevance. By calculating the distance between heatmaps, the Differences between different data points in network decisions are quantified, which helps the model to identify the location of backdoor triggers. The processed heatmaps are categorized using clustering methods to distinguish normal and manipulated data points. Comparing these two clustering methods, Heatmap clustering has a significantly higher computational cost than activation clustering, which is caused by Gromov-Wasserstein to calculate the distances between heatmaps. However, it is more adaptable and robust to different attack patterns and data characteristics.
@@ 4.3.2 Model Identification and Construction
Figure 10: The workflow of meta-classifier defense [131].
Reconstructing model: Model reconstruction methods directly target suspicious models to eliminate malicious backdoor. Liu et al. [146] propose a defense method called “Fine-Pruning”, which consists of two methods: pruning and fine-tuning. The principle of utilizing the pruning is to remove neurons that are inactive under normal input data, for these neurons may be utilized by attackers to implant a backdoor. Fine-tuning is based on the principle of tuning the pre-trained model, which depends on adjusting the weights of the local model to mitigate the impact on backdoor attacks. Combining these two approaches, the strategy in this paper can reconstruct the model to effectively resist the backdoor attack. It can not only maintains the accuracy against clean inputs, but also significantly reduces the success rate of the backdoor attack, especially when complex pruning-aware attacks occur. To purify deep neural networks contaminated by backdoor attacks, Wu et al. [147] proposed a model reconstructing method of Adversarial Neuron Pruning. The core idea is to add perturbations to neurons and prune sensitive neurons to eliminate backdoor attacks. This method can reshape the model based on a small amount of clean data, and has wide applications in resisting backdoor attacks.
Reliable model framework: Zhou et al. [148] built a training framework upon the concept of split learning, designed to prevent backdoor attacks on semantic symbols by ensuring secure data transmission between the sender and receiver. Its U-shaped structure facilitates bidirectional data transmission: forward propagation begins at the sender, continues through a wireless channel to the receiver, where an activation signal is generated at the (k + m)th layer and returned to the sender. The sender then uses this signal to complete forward propagation and compute gradients, which are ultimately sent back to the receiver to update the model parameters. The purpose of this framework is to prevent backdoor attackers from injecting triggers into training data to implant backdoors within the model. By assuming an ideal channel, this framework effectively reduces the risk of data tampering during transmission, thereby preventing backdoor attacks and safeguarding the security and integrity of the SemCom system.
Lesson 3
@@ 4.4 Adversarial Training
Figure 11: Process of adversarial training [149].
@@ 4.4.1 Adversarial Sample Generation
Figure 12: Semantic distance minimization mechanism [152].
@@
4.4.2 Model Optimization
Adversarial training enhances a model’s resilience by teaching it to identify both regular data and adversarial samples—intentionally crafted inputs meant to deceive the model. By iteratively improving its ability to counter these attacks, the model gains robustness against real-world noise and malicious interference. This process requires balancing performance on standard and adversarial data while refining the model’s decision boundaries to bolster generalization and defense. The discussion focuses on three common approaches including adding adversarial regularization, using adaptive disturbance constraint and matching pluggable modules.
Adding adversarial regularization: Adversarial regularization involves adding terms to the loss function, including class consistency, semantic consistency, or distribution consistency between normal and classified adversarial samples, to improve the robustness of adversarial training.
  • 1.
@@ 4.4.2 Model Optimization
Adversarial training with adaptive disturbance constraint: Parameter configuration is crucial to enhance the robustness of the model. In most cases, the parameters of the attack including the number of iterations K, the attack step size $\alpha$ with the perturbation constraint $\varepsilon$, which are artificially predefined during training. However, some researchers argue that individual data points may have different intrinsic robustness, which means they have different distances to the decision boundary of the classifier. If a fixed perturbation tolerance is imposed on all samples at the decision boundary, samples close to the decision boundary will be misclassified when the perturbation value is relatively large. The model is forced to adjust the decision boundary in order to adapt to these perturbed samples, resulting in a distortion of the decision surface of the entire model, which affects the model’s classification performance for other samples, especially samples far from the decision boundary. However, traditional adversarial training uses a uniform treatment for all samples. To address this problem, a series of adaptive methods can be used when setting such parameters, which can better improve the model robustness.
  • 1.
@@
Matching pluggable modules: Pluggable modules provide significant advantages for secure SemCom systems, enhancing security while preserving both efficiency and flexibility. Due to their high compatibility with existing architectures, these modules enable seamless integration without requiring modifications to the existing system structure. He et al. [159] investigate the advantages of secure-aware SemCom systems in countering adversarial attacks. They proposes installing a matched pair of pluggable modules—one positioned after the semantic transmitter and another before the semantic receiver. The transmitter module employs a trainable Adversarial Residual Network to generate adversarial examples, while the receiver module uses a corresponding trainable network to mitigate adversarial attacks and channel noise. To counter semantic eavesdropping, these trainable network are jointly optimized to minimize a weighted sum of the adversarial attack’s strength, mean squared error in SemCom, and the eavesdropper’s confidence in accurately retrieving private information. Numerical results demonstrate that this approach effectively misleads eavesdroppers while preserving high-quality SemCom.
Lesson 4
Adversarial training, which involves adding adversarial attacks to training samples, is favored by researchers for enhancing the robustness of models in SemCom by enabling them to recognize semantic adversarial samples generated through noise addition. Adversarial training approaches can be divided into adversarial sample generation and model optimization. The former includes iterative fast gradient symbol method [151], semantic distance minimization mechanism [152], and hybrid adversarial training strategy [153]. The latter includes adding adversarial regularization [154, 155, 156], adversarial training with adaptive disturbance constraint [157, 158], and matching pluggable modules [159].
@@ 4.5 Differential Privacy
Figure 13: Differential privacy mechanism [161].
@@
Figure 14: The three phase of differential privacy deployed in deep learning model [162].
@@ 4.5.1 Parameter Value Protection
Weight of Model: In AI-enabled networks, differential privacy is widely applied to FL. In FL, a client, including a cell phone or edge device, trains a model based on its local data and uploads updated model weights to a central server. These parameters may leak the client’s data characteristics and thus need to be noised to protect privacy. Differential privacy enables the model updates without the risk of leaking too much information about any individual data point in the training data set [163]. For instance, Wei et al. [164] added random Gaussian noise at the aggregated global model to hide any single client’s update in the centralized pattern. Specifically, Kang Wei et al. propose a new framework based on the DP concept, Noising before Aggregation FL, to effectively prevent information leakage. Each client locally perturbs its training parameters by purposefully adding noise to ensure that the uploaded parameters satisfy a specific level of privacy protection and uploads the updated model weights to the central server before uploading its training parameters to the server for aggregation. The system achieves an effective balance between privacy protection and model performance by sharing key parameters of the model to improve learning accuracy while protecting the privacy of participants’ data. In SemCom systems, the sender can add the noise to the model update before model-sharing, which is helpful for privacy-preserving, especially in multi-hop scenarios. Several researchers [164, 85, 165, 166] have demonstrated that the obfuscated parameters in differential privacy lead to a trade-off between the model training-accuracy and security. Therefore, adding well-designed differential privacy and the proper way to add noise matters to the model-driven networks, especially in the multi-hop manners.
@@
Gradient of Model: The gradient leakage data attack can reconstruct the local data through the shared gradients of the participants, which brings a new challenge to the privacy preservation problem of the learning model. To better address the model security issue, differential privacy is introduced in optimization algorithms for gradient. Specifically, noise is injected into the gradient of the client parameters before the training results are sent to the server. Differential privacy ensures that no information about any individual training sample can be accurately inferred even when the model parameters are made public. A prime example of this is the selective sharing of model parameters during model training by Shokri et al. [163]. The parameter selection rate determines the number of parameters updated in iterations for each parameter subset. During the parameter subset selection process, the gradients whose median values are larger than a specific threshold are randomly selected for uploading, and this approach is similar to the idea of differential privacy. Also, to further protect privacy, differential privacy techniques are used through the Lapp system to ensure that parameter updates do not reveal too much information about individual points in the training data. This is accomplished through sparse vector techniques and a Laplacian mechanism that adds noise to selected gradients and controls the privacy budget. Song et al. [167] introduced differential privacy in Stochastic gradient descent. They theoretically derived the gradient descent with differentially private algorithm and verified the feasibility of the algorithm. Abadi et al. [168] developed the algorithm on this basis and proposed the differentially private stochastic gradient descent algorithm (DP-SGD) algorithm, which is reflected in the proposal of an effective algorithm for calculating the single training examples’ gradient, the efficient algorithm for subdividing the task into smaller batches to reduce memory footprint, and the application of the differentially private principal projection at the input layer. Specifically, the algorithm adds Gaussian noise to the gradient to preserve privacy, where the size of the noise is proportional to the sensitivity of the gradient. It controls the impact of each training sample on the model update by limiting the l2 norm of the gradient through Norm clipping, and uses the “Moments Accountant algorithm to compute the overall privacy loss during the training process to measure whether a set level of privacy protection is met. Liu et al. [169] introduces the Adaptive Privacy Budget-based Differential Privacy (APB-DP) method. APB-DP leverages differential privacy (DP) with an adaptive privacy budget, adjusting noise levels as models converge to ensure consistent privacy protection throughout training. Additionally, APB-DP accounts for wireless channel effects to reduce interference.Experimental results demonstrate that APB-DP decreases privacy leakage by 13 percent and minimizes performance loss by 71 percent compared to leading DP-based training methods, providing a more secure and efficient solution for SemCom.

### v2

@@ 4.3 Defensive Methods of Backdoor Attacks
The below, as shown in Table LABEL:backdoorattacks, the following content, based on multiple academic studies, delves into various defensive methods of backdoor attacks in different stages and their potential applications in SemCom.
Table 7: List of Representative Backdoor Defense Schemes
[…]
Figure 9: Classification of neural network backdoor defense [139], where the process of data collection and model training is referred to as model training, while the stages of model testing and model deployment are considered subsequent to model training. The phases of training and post-training correspond to two distinct types of neural network backdoor attacks: one in which the user acquires training data from a third party, and the other in which the user obtains a model from a third party.
@@ 4.3.1 Data Processing
    Clustering based Defense. For contaminated datasets, toxic data can be separated from normal data according to cluster analysis. Activation Clustering is a defense method proposed by Chen [133] et al. The key idea is to analyze the activation mode of network for the training data, to determine whether the data is poisoned and which data points are toxic. For each sample in the training set, AC records the activation output of the hidden layer. These activation outputs are flattened into one-dimensional vectors and classified based on their labels. After dimensionality reduction, these vectors are clustered and analyzed into toxic and normal patterns. The data is analyzed by exclusionary reclassification, silhouette score, size comparison and many other ways to filter out the toxic data. Based on the labeled toxic data, the system can further train the model until convergence to achieve the purpose of backdoor repairing. One more clustering approach is Heatmap Clustering proposed by Schulth et al. [141]. This method utilizes Explainable Artificial Intelligence techniques, specifically Layer-wise Relevance Propagation, to generate heatmaps. The heatmap shows the correlation between each pixel point of the input sample and the network predictions, which can be visualized as a two-dimensional array where the color shades indicate the strength of the relevance. By calculating the distance between heatmaps, the Differences between different data points in network decisions are quantified, which helps the model to identify the location of backdoor triggers. The processed heatmaps are categorized using clustering methods to distinguish normal and manipulated data points. Comparing these two clustering methods, Heatmap clustering has a significantly higher computational cost than activation clustering, which is caused by Gromov-Wasserstein to calculate the distances between heatmaps. However, it is more adaptable and robust to different attack patterns and data characteristics.
@@ 4.3.2 Model Identification and Construction
Figure 10: The workflow of meta-classifier defense [139], where the data that reflect the characteristics of both benign and backdoored models are extracted firstly; this data includes key information such as the models’ internal parameters and output behaviors. Subsequently, these features are utilized to train a meta-classifier, which is a supervised learning model capable of learning to distinguish between benign and backdoored models based on these features. Once trained, the meta-classifier is employed to predict the nature of new model samples, determining whether they are benign or backdoored, with 0 representing a benign model and 1 representing a backdoored model. The entire process forms a closed loop, commencing with feature extraction, proceeding through training and prediction, and ultimately applying to the security detection of the target model. In the event that a backdoored model is detected, the system can implement corresponding measures, such as rejection, retraining, or model repair.
Reconstructing model: Model reconstruction methods directly target suspicious models to eliminate malicious backdoor. Liu et al. [137] propose a defense method called “Fine-Pruning”, which consists of two methods: pruning and fine-tuning. The principle of utilizing the pruning is to remove neurons that are inactive under normal input data, for these neurons may be utilized by attackers to implant a backdoor. Fine-tuning is based on the principle of tuning the pre-trained model, which depends on adjusting the weights of the local model to mitigate the impact on backdoor attacks. Combining these two approaches, the strategy in this paper can reconstruct the model to effectively resist the backdoor attack. It can not only maintains the accuracy against clean inputs, but also significantly reduces the success rate of the backdoor attack, especially when complex pruning-aware attacks occur. To purify DNNs contaminated by backdoor attacks, Wu et al. [149] proposed a model reconstructing method of Adversarial Neuron Pruning. The core idea is to add perturbations to neurons and prune sensitive neurons to eliminate backdoor attacks. This method can reshape the model based on a small amount of clean data, and has wide applications in resisting backdoor attacks.
Reliable model framework in Semcom: Zhou et al. [138] explored backdoor attacks that specifically target deep learning-based SemCom systems. Given that current studies on backdoor attacks are not adapted to the unique context of SemCom, a novel backdoor attack model, referred to as backdoor attack on semantic symbols, is introduced. Based on this model, appropriate countermeasures are proposed. In particular, a training framework is developed to mitigate the risks of Backdoor Attack on Semantic Symbols. Furthermore, defense strategies utilizing reverse engineering and pruning techniques are formulated to safeguard SemCom systems from backdoor attacks. Simulation results validate the effectiveness of both the proposed attack model and the corresponding defense mechanisms.
Lesson 2
@@ 4.4 Adversarial Training
Figure 11: Process of adversarial training [150], where adversarial training enhances traditional training methods by incorporating additional steps into the training process, including the generation of adversarial data and the prediction of results.Initially, data is fed into the model alongside their corresponding labels to produce prediction results, which are then compared against the true labels to calculate the loss. This loss is subsequently used to perform backpropagation, adjusting the model’s parameters to improve its accuracy on the given data. Additionally, the model undergoes an extra step where it is trained with adversarial data, which has been intentionally modified to challenge the model’s robustness. This adversarial data is generated through adversarial attacks designed to deceive or mislead the model. The model’s performance on both standard and adversarial data is assessed, with the prediction results of adversarial data being a critical component in evaluating the model’s resilience against such attacks. This dual training process aims to enhance the model’s ability to generalize well and maintain accuracy under both normal and adversarial conditions, thereby strengthening its overall robustness against potential threats in real-world applications.
The below, as shown in Table LABEL:adversarialtraining, the following content, based on multiple academic studies, delves into various adversarial training methods for data level and model level, and their potential applications in SemCom.
Table 8: List of Representative Adversarial Training Schemes
[…]
@@ 4.4.1 Adversarial Sample Generation
Figure 12: Semantic distance minimization mechanism [152], where lines of the same color represent the same data stream, while the dashed line indicates that the semantic encoder is updated based on the loss function.The diagram outlines a four-step process for generating and utilizing adversarial data to enhance a model’s robustness. In step 1, adversarial data is produced by leveraging a knowledge base and benign data to magnify the semantic distance, creating attack data intended to challenge the model. In step 2, a semantic encoder is employed to extract semantic representations from both the benign and adversarial data. Step 3 involves calculating the loss, which combines the natural loss of the model with a Kullback-Leibler (KL) divergence term, weighted by a factor $\lambda$, to quantify the difference between the semantic representations of the benign and adversarial data. Step 4 aims to minimize this semantic distance, thereby reducing the disparity between the representations and enhancing the model’s ability to generalize across both normal and adversarial conditions. This process is crucial for improving the model’s resilience against potential semantic attacks and ensuring its reliability in various scenarios.
Defense Framework in multi-user Semcom: Peng et al. [90] proposes an effective attack-defense game framework, specifically designed to defend against data poisoning attacks during image transmission. The proposed method used reinforcement learning to determine the optimal attack strategy for each attack type, aiming to strengthen the attack while evading detection. To generate corresponding adversarial samples, we design an Adversarial Sample Generator based on Conditional Generative Adversarial Networks. Next, an Attack Defender is introduced to detect data poisoning attacks, exclude the poisoned samples from the learning process of the target model, and generate confrontation samples guided by the optimal attack strategy to enhance the robustness of the defender. Simulation results demonstrate that the framework successfully identifies the optimal attack strategy, which further degrades the target model’s accuracy while maintaining a high evasion rate. Additionally, the generator produces effective adversarial samples, and the defender outperforms five state-of-the-art detectors in terms of Top-1 accuracy on three widely used image datasets, tested under the AWGN channel.
@@
4.4.2 Adversarial Regularization
Adversarial training enhances a model’s resilience by teaching it to identify both regular data and adversarial samples—intentionally crafted inputs meant to deceive the model. By iteratively improving its ability to counter these attacks, the model gains robustness against real-world noise and malicious interference. This process requires balancing performance on standard and adversarial data while refining the model’s decision boundaries to bolster generalization and defense.
Adversarial regularization is a technique used to improve the robustness of machine learning models, particularly in the context of adversarial training. It works by incorporating additional terms into the loss function that help the model to better defense adversarial attacks, including class consistency, semantic consistency, or distribution consistency. These terms typically focus on ensuring consistency between normal samples and adversarial examples in various dimensions to improve the robustness of adversarial training.
@@ 4.4.2 Adversarial Regularization
4.4.3 Adaptive Disturbance Constraint
Parameter configuration plays a critical role in enhancing the robustness of a model. Typically, the parameters of the attack including the number of iterations K, the attack step size $\alpha$ with the perturbation constraint $\varepsilon$, which are artificially predefined during training. However, some researchers argue that individual data points may have different intrinsic robustness, which means they have different distances to the decision boundary of the classifier. If a fixed perturbation tolerance is imposed on all samples at the decision boundary, samples close to the decision boundary will be misclassified when the perturbation value is relatively large. The model is forced to adjust the decision boundary in order to adapt to these perturbed samples, resulting in a distortion of the decision surface of the entire model, which affects the model’s classification performance for other samples, especially samples far from the decision boundary. However, traditional adversarial training uses a uniform treatment for all samples. To address this problem, a series of adaptive methods can be used when setting such parameters, which can better improve the model robustness.
@@ 4.4.3 Adaptive Disturbance Constraint
4.4.4 Matching Pluggable Modules
Pluggable modules offer substantial benefits for SemCom systems by significantly enhancing security while maintaining both efficiency and flexibility. These modules are designed with high compatibility, allowing them to be easily integrated into existing system architectures without the need for extensive modifications. By providing such integration, pluggable modules ensure that security enhancements can be implemented with minimal disruption to the overall system performance. This adaptability not only helps in safeguarding sensitive communication but also preserves the system’s operational efficiency, making it an ideal solution for evolving communication environments where flexibility is essential.
He et al. [158] investigate the advantages of secure-aware SemCom systems in countering adversarial attacks. They proposes installing a matched pair of pluggable modules—one positioned after the semantic transmitter and another before the semantic receiver. The transmitter module employs a trainable Adversarial Residual Network to generate adversarial examples, while the receiver module uses a corresponding trainable network to mitigate adversarial attacks and channel noise. To counter semantic eavesdropping, these trainable network are jointly optimized to minimize a weighted sum of the adversarial attack’s strength, mean squared error in SemCom, and the eavesdropper’s confidence in accurately retrieving private information. Numerical results demonstrate that this approach effectively misleads eavesdroppers while preserving high-quality SemCom.
Lesson 3
Adversarial training, which involves adding adversarial attacks to training samples, is favored by researchers for enhancing the robustness of models in SemCom by enabling them to recognize semantic adversarial samples generated through noise addition. Adversarial training approaches can be divided into adversarial sample generation, adversarial regularization, adaptive disturbance constraint, and matching pluggable modules. The adversarial sample generation schemes includes iterative fast gradient symbol method [151], semantic distance minimization mechanism [152], and hybrid adversarial training strategy [153]. The adversarial regularization includes combination adversarial training and encryption [154], application for task-oriented SemCom [155], and trade off between the accuracy and robustness [160]. The adaptive disturbance constraint methods are applied in max-margin adversarial training [156] and Semcom [157]. Matching pluggable module is uesd in SemCom system to defense adversarial attacks [158].
@@ 4.5 Differential Privacy
Figure 13: Differential privacy mechanism [162], where depicts a privacy-preserving system architecture where an Analyst submits a query to a Privacy Guard, which acts as an intermediary between the Analyst and the Database. The Privacy Guard is responsible for processing the query in a way that protects sensitive information within the database. It does this by adding noise to the query results before sending them back to the Analyst, thus providing a Noisy Response. This mechanism ensures that while the Analyst can still receive useful information, the privacy of the data in the Database is maintained by preventing the Analyst from gaining access to exact or potentially sensitive data points.
@@
Figure 14: The three phase of differential privacy deployed in DL model [163], where the phase of the mechanism is divided it into three types: input layer, hidden layer, and output layer. Firstly, sensitive data undergoes differential privacy training to generate synthetic data, ensuring that the privacy of individuals within the dataset is maintained. This synthetic data is then used for analysis within a neural network composed of an input layer, one or more hidden layers, and an output layer. To further enhance privacy, noise is injected into the gradients during the training process, and the objective function can be perturbed. These measures are designed to prevent the reconstruction of the original data from the model, thereby safeguarding sensitive information.
The below, as shown in Table LABEL:differentialprivacy, the following content, based on multiple academic studies, delves into various differential privacy methods in different stages and their potential applications in SemCom.
Table 9: List of Representative Differential Privacy Schemes
[…]
@@ 4.5.1 Parameter Value Protection
Weight of Model Protection: Wei et al. [164] added random Gaussian noise at the aggregated global model to hide any single client’s update in the centralized pattern. Specifically, Kang Wei et al. propose a new framework based on differential privacy, Noising before Aggregation FL, to effectively prevent information leakage. Each client locally perturbs its training parameters by purposefully adding noise to ensure that the uploaded parameters satisfy a specific level of privacy protection and uploads the updated model weights to the central server before uploading its training parameters to the server for aggregation. The system achieves an effective balance between privacy protection and model performance by sharing key parameters of the model to improve learning accuracy while protecting the privacy of participants’ data.
@@
Gradient of Model Protection: Shokri et al. [165] chose to share the model parameter selectively. The parameter selection rate determines the number of parameters updated in iterations for each parameter subset. During the parameter subset selection process, the gradients whose median values are larger than a specific threshold are randomly selected for uploading, and this approach is similar to the idea of differential privacy. Song et al. [171] introduced differential privacy in Stochastic gradient descent. They theoretically derived the gradient descent with differentially private algorithm and verified the feasibility of the algorithm. Abadi et al. [172] developed the algorithm on this basis and proposed the differentially private stochastic gradient descent algorithm, which is reflected in the proposal of an effective algorithm for calculating the single training examples’ gradient, the efficient algorithm for subdividing the task into smaller batches to reduce memory footprint, and the application of the differentially private principal projection at the input layer. Specifically, the algorithm adds Gaussian noise to the gradient to preserve privacy, where the size of the noise is proportional to the sensitivity of the gradient.
Parameter Protection in Semcom: Liu et al. [166] introduces the Adaptive Privacy Budget-based Differential Privacy (APB-DP) method. APB-DP leverages differential privacy with an adaptive privacy budget, adjusting noise levels as models converge to ensure consistent privacy protection throughout training. Additionally, APB-DP accounts for wireless channel effects to reduce interference. Experimental results demonstrate that APB-DP decreases privacy leakage by 13 percent and minimizes performance loss by 71 percent compared to leading differential privacy-based training methods, providing a more secure and efficient solution for SemCom. In SemCom systems, the sender can add the noise to the model update before model-sharing, which is helpful for privacy-preserving, especially in multi-hop scenarios. Several researchers [164, 85, 173, 174] have demonstrated that the obfuscated parameters in differential privacy lead to a trade-off between the model training-accuracy and security. Therefore, adding well-designed differential privacy and the proper way to add noise matters to the model-driven networks, especially in the multi-hop manners.